OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. 9|9 – 8x| = 2x + 3

OpenStudy (anonymous):

solve two equations \[9(9-8x)=2x+3\] and \[9(9-8x)=-2x-3\]

OpenStudy (anonymous):

just curious before we start is that first 9 for real, or is it a typo?

OpenStudy (anonymous):

forreal

OpenStudy (anonymous):

ok then first one gives \[81-72x=2x+3\] \[81=74x+3\] \[78=81x\] \[x=\frac{78}{81}\] are you sure? in any case second one is similar

OpenStudy (anonymous):

yea im sure it just didnt seem right when i did it. so i wanted to see how other people did it..